If f(x)= log(2,x) and g(x)= 2x^2 + 14, determine the value of (f o g) (5)?
The "o" in the (f o g) is a symbol that I was not able to do, and I don't think it represents multiplication...or does it? Here is how I thought it would be done:
f(g(5)) = log (base 2)(2(5^2) + 14)

= log (2, 64)

= 2^x =64
x = 6
Is this right?

Answer by stanbon(75887)

If f(x)= log(2,x) and g(x)= 2x^2 + 14, determine the value of (f o g) (5)?
The "o" in the (f o g) is a symbol that I was not able to do, and I don't think it represents multiplication...or does it? Here is how I thought it would be done:
f(g(5)) = log (base 2)(2(5^2) + 14)
= log (2, 64)
= 2^x =64
x = 6
Is this right?
----------------------------------
fog(x) means f[g(x)]
----
Think of f being a fish and g being a guppy(minnow).
You feed the guppy a grain of food (x)
Then the fish eats the guppy.
----------------------------------------
f(x)= log(2,x) and g(x)= 2x^2 + 14, determine the value of (f o g) (5)
fog(5) = f[g(5)] = f[2*5^2+14] = f[64] = log(base2)64 = 6
================================================================
You are correct.
================================================================
Cheers,
Stan H.